C and Y candy company mixes candy that costs $6.00 per kg with candy that costs $4.50 per kg. How many kilograms of each are needed to make ao 3kg box that costs $15.00?
Answer by mananth(16949)

$ 6 Candy x kgs
$4.5 candy y kgs
x+y=3....................(1)
6x+4.5y=15................(2)
multiply (1) by -6
-6x-6y=-18
add to (2)
-1.5y=-3
/-1.5
y=2
x=1
$ 6 Candy 1 kg
$4.5 candy 2 kgs